OpenStudy (briellshaw):

Vince worked 462 hours in 11 weeks. At what rate did he work in hours per week? 41 hours per week 44 hours per week 43 hours per week 42 hours per week

OpenStudy (anonymous):

hours per week is the same as hours/week

OpenStudy (briellshaw):

462/11=?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\(\sf\Large\frac{462}{11}=?\)

OpenStudy (ohohaye):

yep

OpenStudy (anonymous):

Use a calculator and divide the fraction.

OpenStudy (briellshaw):

\[43/11\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Huh? you had it right the first time

OpenStudy (briellshaw):

sorrry \[462/11\]

OpenStudy (anonymous):

Go on, and divide that.

OpenStudy (briellshaw):

\[462/11=42\]

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Good job @BriellShaw \(\Large\ddot\smile\)

OpenStudy (briellshaw):

next question

OpenStudy (briellshaw):

Sam is filling up his pool. The water hose filled the 336 gallon pool in 48 minutes. At what rate did the hose fill the pool in gallons per minute? 8 gallons per minute 6 gallons per minute 7 gallons per minute 9 gallons per minute

OpenStudy (anonymous):

Make a new post.